The square of 9 more than a number is equal to 9 more than the square of the number. What is the number?
thanx

Answer by fractalier(6550)

Let the number be x. Then we have
(x + 9)^2 = x^2 + 9
x^2 + 18x + 81 = x^2 + 9
18x + 81 = 9
18x = -72
x = -4